xx FishboneMoncriefID: Black Hole Accretion Disk Initial Data

FishboneMoncriefID: Black Hole Accretion Disk Initial Data

Author: Zachariah B. Etienne Date: 2019-09-28 (Sat, 28 Sep 2019)

This thorn is documented within two Jupyter notebooks, contained within the FishboneMoncriefID/doc/ directory:

Tutorial-FishboneMoncriefID.ipynb: Contains basic description, equations, and citations for this initial data type.

     Tutorial-ETK_thorn-FishboneMoncriefID.ipynb: Constructs this Einstein Toolkit thorn.

1 Parameters

11randomxmax 10Scope: privatex01 REAL

11Description: Ceiling value of random perturbation to initial pressure, where perturbed pressure = pressure*(1.0 + (randomxmin + (randomxmax-randomxmin)*RAND[0,1)))

11 Range 10 x01 Default: 0.02
11*:*  10Any value

11

11randomxmin  10Scope: privatex01  REAL

11Description: Floor value of random perturbation to initial pressure, where perturbed pressure = pressure*(1.0 + (randomxmin + (randomxmax-randomxmin)*RAND[0,1)))

11 Range 10 x01 Default: -0.02
11*:*  10Any value

11a  10Scope: restrictedx01  REAL

11Description: The spin parameter of the black hole

11 Range 10 x01 Default: 0.9375
110:1.0  10Positive values, up to 1. Negative disallowed, as certain roots are chosen in the hydro fields setup. Check those before enabling negative spins!

11

11axb  10Scope: restrictedx01  REAL

11Description: Scaling factor for the vector potential

11  Range  10  x01  Default: 1.0

11*:*  10

11

11gamma  10Scope: restrictedx01  REAL

11Description: Equation of state: P = kappa * rhoxamma

11 Range 10 x01 Default: 1.3333333333333333333333333333
110.0:*  10Positive values

11

11kappa  10Scope: restrictedx01  REAL

11Description: Equation of state: P = kappa * rhoxamma

11 Range 10 x01 Default: 1.0e-3
110.0:*  10Positive values

11

11m  10Scope: restrictedx01  REAL

11Description: Kerr-Schild BH mass. Probably should always set M=1.

11 Range 10 x01 Default: 1.0
110.0:*  10Must be positive

11

11rxatxmaxxdensity 10Scope: restrictedx01  REAL

11Description: Radius at maximum disk density. Needs to be > rxin

11 Range 10 x01 Default: 12.0
110.0:*  10Must be positive

11

11rxin  10Scope: restrictedx01  REAL

11Description: Fixes the inner edge of the disk

11 Range 10 x01 Default: 6.0
110.0:*  10Must be positive

11

11initialxhydro  10Scope: shared from HYDROBASEx01  KEYWORD

11Extends ranges:

11FishboneMoncriefID 10Initial GRHD data from FishboneMoncriefID solution

11

2 Interfaces General Implements: fishbonemoncriefid

Inherits: admbase grid hydrobase

3 Schedule This section lists all the variables which are assigned storage by thorn WVUThorns/FishboneMoncriefID. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11ADMBase::metric[metricxtimelevels] ADMBase::curv[metricxtimelevels] ADMBase::lapse[lapsextimelevels] ADMBase::shift[shiftxtimelevels]11x

11x  11x

11

Scheduled Functions

HydroBasexInitial x fishbonemoncriefxetxgrhdxinitial x set up general relativistic hydrodynamic (grhd) fields for fishbone-moncrief disk

11x11Language:11c
11 x 11Reads: 11grid::x(everywhere)
11 x 11x  11grid::y(everywhere)
11 x 11x  11grid::z(everywhere)
11 x 11Type: 11function
11 x 11Writes: 11admbase::alp(everywhere)
11 x 11x  11admbase::betax(everywhere)
11 x 11x  11admbase::betay(everywhere)
11 x 11x  11admbase::betaz(everywhere)
11 x 11x  11admbase::kxx(everywhere)
11 x 11x  11admbase::kxy(everywhere)
11 x 11x  11admbase::kxz(everywhere)
11 x 11x  11admbase::kyy(everywhere)
11 x 11x  11admbase::kyz(everywhere)
11 x 11x  11admbase::kzz(everywhere)
11 x 11x  11admbase::gxx(everywhere)
11 x 11x  11admbase::gxy(everywhere)
11 x 11x  11admbase::gxz(everywhere)
11 x 11x  11admbase::gyy(everywhere)
11 x 11x  11admbase::gyz(everywhere)
11 x 11x  11admbase::gzz(everywhere)
11 x 11x  11hydrobase::vel(everywhere)
11 x 11x  11hydrobase::rho(everywhere)
11 x 11x  11hydrobase::eps(everywhere)
11 x 11x  11hydrobase::press(everywhere)

11

CCTKxINITIAL x fishbonemoncriefxetxgrhdxinitialxxperturbxpressure x add random perturbation to initial pressure, after seed magnetic fields have been set up (in case we’d like the seed magnetic fields to depend on the pristine pressures) x

11x11After: 11seedxmagneticxfields
11 x 11Before: 11illinoisgrmhdxidxconverter
11 x 11Language:11c
11 x 11Type: 11function